OpenStudy (anonymous):

Yuma used 1,040 ft of fence to enclose a rectangular pasture. One side borders a river where there is already a thick hedge. That side needs no fencing. c. Write an expression for the area of the pasture in terms of x only. d. Suppose Yuma wants the pasture to enclose at least 60,000 square feet. Write a sentence relating your answer in Part c to the area the fence must enclose.

OpenStudy (anonymous):

I need help with part c!

zepdrix (zepdrix):

So for part a) did you get,\[\Large 2x+L=1040 \qquad\to\qquad L=1040-2x\]And for b) I guess we would get,\[\Large A=xL\]Right?

OpenStudy (anonymous):

Exactly, but I can't figure out part c

zepdrix (zepdrix):

\[\Large \color{royalblue}{L=1040-2x} \qquad\to\qquad A=x\color{royalblue}{L}\]Plug! :)

OpenStudy (anonymous):

Thank you so much!!! It seems so obvious now!

zepdrix (zepdrix):

heh :3

OpenStudy (anonymous):

Ah!!! No, I meant part d, i had c, but part d?

zepdrix (zepdrix):

Hmm so for d we want the `area` to equal 60,000.\[\Large \color{royalblue}{A=60000} \qquad\to\qquad \color{royalblue}{A}=x(1040-2x)\]I'm not really sure what they want in your sentence :P hmm

zepdrix (zepdrix):

Maybe say something like... When we let our A=60000, it allows us to solve for one of our dimensions, x.